OpenStudy (anonymous):

23 = 7x HELP PLEASEEEEE!!!

OpenStudy (anonymous):

simple equation just divide by 7 on both sides: 23 = 7x /7 /7 x = ?

OpenStudy (anonymous):

3.28

OpenStudy (anonymous):

I will post the answers hold on 1 minute lol

OpenStudy (anonymous):

ok i thought so

OpenStudy (anonymous):

A) X = 23/7 B) X = 7/23 C) 3 1/7 D) X = 3

OpenStudy (anonymous):

i became your fan@lynfan

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

the answer is a) based off of my description

OpenStudy (anonymous):

yeah I figured @twistnflip i picked 3 and got it wrong they said it was A